# Analysis and Optimized CXL-Attached Memory Allocation for Long-Context LLM Fine-Tuning

### v1

@@ II-C CXL-Attached Memory
For CXL-attached memory, Figure 4 illustrates the differences in data paths and latency between local and Type 3 devices. Accessing local memory follows a direct path from the CPU cores through the CPU cache and memory controller, resulting in latencies between 80 and 140 nanoseconds [27]. In contrast, accessing CXL-attached memory involves traversing the PCIe interface using the CXL.io and CXL.mem protocols. This path requires coordination between the CPU and the CXL memory controllers, leading to increased latency ranging from 170 to 250 nanoseconds [27]. While CXL-attached memory is utilized as a system memory extension, the Linux kernel integrates them as Non-Uniform Memory Access (NUMA) nodes [35]. This integration allows the system to manage CXL-attached memory within the existing memory management framework. Without manual intervention, the kernel uses NUMA policies to determine how memory is allocated across nodes. However, users can manually control memory allocation using tools like numactl and libnuma [43]. For example, a naive interleave policy distributes memory pages in a round-robin fashion among available NUMA nodes [44]. Although existing integration effectively enlarges usable system memory, our investigation reveals that the performance of CPU-offloaded long-context LLM is suboptimal (See Section III).
Figure 5: Latency comparison of the CPU-Based Adam optimizer step.
(a) Single GPU
(b) Dual GPU
Figure 6: Bandwidth Comparison of System Memory to GPU Memory Transfers.
(a) Single GPU
(b) Dual GPU
Figure 7: Latency breakdown of CPU offloading of local DRAM vs. CXL memory.
III CXL-Attached Memory Performance Analysis
This section empirically characterizes the performance impact of storing CPU‑offloaded data to CXL‑attached memory. The naive deployment first prolongs CPU‑based optimizer steps (See Section III-A) and induces severe bandwidth contention on the shared PCIe link in multi‑GPU configurations (See Section III-B). These findings underscore the necessity of CXL‑aware data‑placement strategies (See Section III-C)
III-A CPU‑Offloading Slowdowns on CXL-Attached Memory
@@
As illustrated in Section II-A, the long‑context CPU‑offloading workflow stores full‑precision parameters, gradients, and optimizer states in system memory so that the CPU can perform the optimizer update locally. As each parameter update is independent, the optimizer phase exhibits ample parallelism. Practical implementations, such as ZeRO-Offload, exploit OpenMP threads and SIMD instructions (e.g., AVX2) to accelerate this compute-intensive step [20]. Consequently, the optimizer phase is highly parallelized and sensitive to increased latency when accessing offloaded data structures. To quantify how memory placement affects optimizer latency within the long‑context CPU‑offloading workflow, the CPU‑based Adam optimizer is benchmarked with offloaded data structures residing either in local DRAM or in CXL-attached memory. Figure 7 summarizes the results. For each configuration, the data size is varied to emulate different LLM scales, excluding trivially small cases. Hardware details are listed in Table II with Config. A.
@@
In Figure 7, an “element” consists of a 4 bytes parameter, a 4 bytes gradient, and 8 bytes of optimizer state. The SIMD kernel processes each element in three steps: (i) it loads the parameter, gradient, and state from memory into vector registers; (ii) executes the floating‑point update; and (iii) writes the updated values back. For modest data volumes, the latency penalty of CXL‑attached memory is negligible; however, once the element count exceeds roughly 20 million, optimizer time on CXL rises sharply, reaching nearly 4 times the DRAM baseline. The primary cause is the higher access latency of the CXL path (170-250 ns) compared to local DRAM (80-140 ns), as shown earlier in Figure 4. These results indicate that naively placing latency‑critical optimizer data in CXL‑attached memory can severely degrade fine‑tuning performance. Effective CXL deployments for long‑context CPU‑offloading need to keep latency‑sensitive data in low‑latency DRAM and relegate latency‑tolerant data to CXL‑attached memory, respectively.
(a) CXL-aware memory allocation
(b) Multi-AIC striping for GPU
(c) Multi-AIC striping for CPU
Figure 8: Illustrative examples of CXL-aware allocation place latency-sensitive CPU compute data (full-precision $P$, $G$, $O$) in local DRAM and latency-tolerant GPU transfer data ($A$, $P$, $G$) in CXL-attached memory. And two Multi-AIC Striping scenarios, one for GPU-related to mitigate bandwidth contention. Another for striping CPU-computed data (e.g. full-precision $O$) across local DRAM and multiple CXL AICs to leverage aggregate memory capacity and bandwidth.
III-B GPU Data Transfers Slowdown on CXL-attached Memory
@@
The CPU‑offloading workflow described in Section II-A entails not only a CPU‑intensive optimizer step but also frequent, high‑volume data movement between system and GPU memory. Before each layer’s computation, model parameters and checkpointed activations are copied from system memory to the GPU; afterwards, the resulting gradients are copied back. To quantify how the physical location of the system memory influences these transfers, this section measures the GPU copy performance with source (or destination) buffers that reside in local DRAM or CXL-attached memory. For each experiment, page‑aligned host buffers are allocated with posix_memalign, and numactl pins each allocation to the desired NUMA node—local DRAM or a CXL Type 3 AIC. The buffers are then registered with cudaHostRegister, enabling direct DMA over PCIe and avoiding extra copies through CPU caches. Asynchronous cudaMemcpyAsync transfers are issued via CUDA streams and events, while latency and effective bandwidth are recorded. Multi‑GPU tests launch concurrent copy requests to both target GPUs. Hardware details are listed in Table II with Config. A. and PCIe Gen5 ×16 link (128 GB/s bidirectional,  64 GB/s per direction) is utilized.
Results are presented in Figure 7. With a single GPU in Figure 7(a), transfer bandwidth from CXL‑attached memory is virtually identical to that from local DRAM; throughput climbs with request size until it reaches the PCIe limit because page‑locked buffers give both memory types an equivalent DMA path, making the operation interface‑bound. A stark contrast emerges in the dual‑GPU case, shown in Figure 7(b). When both GPUs copy concurrently, the aggregate bandwidth from the CXL AIC drops to roughly 25 GiB/s, which is far below the sum of two local DRAM transfers and well short of the theoretical link rate. The bottleneck is the single PCIe connection between the CXL AIC and the host; competing DMA streams must share this finite link, dramatically throttling each transfer. Local DRAM, attached through the CPU’s memory controllers, avoids such shared‑link contention.
Although these single‑GPU results align with prior reports showing that CXL‑to‑GPU copies can rival local DRAM copies under isolated conditions [39, 28], the multi‑GPU results highlight a critical scaling limitation: simultaneous GPU accesses to a single CXL AIC introduce severe bandwidth contention, potentially negating the capacity benefits of CXL‑attached memory in multi‑GPU fine‑tuning workflows.
III-C Motivation: End‑to‑End Fine‑Tuning Slowdown
@@
Based on the characteristics outlined in previous sections, this section measures how naively incorporating CXL‑attached memory affects end‑to‑end performance during LLM fine‑tuning with CPU offloading. The comparison pits a local DRAM baseline against a configuration that combines local DRAM with CXL memory. Table II details the platform: one or two H100 GPUs fine‑tuning a 12‑billion‑parameter model (4K context, batch size 16) while a single CXL Type 3 AIC contributes 512 GiB of capacity. Memory bindings were controlled with numactl: baseline runs allocated pages to the local DRAM NUMA node, whereas CXL runs naively round-robin interleaved allocations to the CXL NUMA node and the 128 GiB local DRAM NUMA node. Figure 7(a) presents the latency profile in the single‑GPU scenario. The optimizer step (STEP phase) suffers the most because its CPU‑bound loads and stores are acutely sensitive to the higher access latency of CXL memory, confirming the expectation from Section III-A. Phases dominated by GPU transfers (FWD, BWD) exhibit smaller slowdowns; prefetching and asynchronous DMA obscure part of the added latency.
Figure 7(b) shows that a dual‑GPU configuration shifts the bottleneck. As predicted in Section III-B, concurrent traffic from both GPUs saturates the single AIC’s PCIe link, so the FWD and BWD phases degrade markedly, while STEP remains limited primarily by latency. Overall, naive CXL adoption introduces two principal issues: (1) substantial latency inflation during the optimizer step, and (2) severe bandwidth contention for GPU transfers in multi‑GPU setups. These findings highlight the need for CXL‑aware data‑placement schemes that position parameters, gradients, optimizer states, and activations according to their sensitivity to latency versus bandwidth, as well as mechanisms that alleviate bandwidth contention across multiple GPUs. Properly addressing these challenges would let systems with modest local DRAM and ample CXL capacity fine‑tune larger models and longer contexts without incurring prohibitive slowdowns.
@@ IV CXL-aware Long-Context LLM Fine-Tuning
To curb the latency hit of CXL while still exploiting its capacity, the proposed CXL-aware long-context LLM fine-tuning first introduces the CXL‑aware memory allocation to keep latency‑critical optimizer data in DRAM and relegates bulky, latency‑tolerant tensors to CXL memory (See Section IV-A). Complementing this, Multi‑AIC striping spreads each GPU’s CXL‑resident data across multiple add‑in cards, eliminating single-link contention and pooling bandwidth, an approach that also scales when optimizer states spill beyond DRAM (See Section IV-B).
IV-A CXL-aware Memory Allocation
Building on the previous observations in Section III-A, the CXL‑aware memory allocation is introduced to reserve local DRAM for the data most sensitive to latency during the CPU‑based optimizer step, namely the full‑precision model parameters $P$, gradients $G$, and optimizer states $O$. Data that tolerates higher latency, such as checkpointed activations $A$ and bf16 copies of parameters and gradients streamed to the GPU, is directed to CXL‑attached memory. This selective placement harnesses the greater capacity of CXL memory for storing latency‑tolerant information without suffering the latency penalty of CXL-attached memory when latency‑critical optimizer data reside on slower medium. Aligning each data type with the memory tier that best matches its access pattern not only improves efficiency but also enhances scalability for long‑context workloads. As the context length or batch size expands, the footprint of the activation checkpoints $A$ increases substantially (see Section II-B), while the sizes of $P$, $G$, and $O$ remain fixed relative to the model parameters. Offloading the ever‑increasing activation data to the expandable CXL pool alleviates pressure on local DRAM, enabling a stable DRAM configuration while relying on CXL for capacity growth. Figure 8(a) visualizes the memory layout, highlighting how each data component is distributed between local DRAM and CXL‑attached memory.
IV-B Multi-AIC Striping
Even though CXL AICs make it straightforward to expand system memory capacity, effectively using multiple cards is still challenging. Earlier observations in Section III-B reveal severe bandwidth contention whenever several GPUs tried to access the same AIC. Additional profiling also showed that latency‑tolerant transfer data—checkpointed activations $A$, parameters $P$, and gradients $G$ benefit most from placement in CXL memory. Taken together, these findings motivate the multi‑AIC striping approach, which splits each GPU process’s data across all available AICs, preventing concurrent traffic from piling onto a single card and relieving the observed bandwidth bottleneck. When the system contains more AICs than GPUs—for instance, one GPU paired with two AICs—the data is striped across all cards so transfers can harness their combined bandwidth. Figure 8(b) illustrates the arrangement for a two‑GPU, two‑AIC configuration.
Multi‑AIC striping extends naturally to CPU‑intensive phases. If the full‑precision optimizer states $O$ exceed local DRAM capacity, the data can be partitioned across both DRAM and the available AICs. During the optimizer step, the CPU accesses these partitions in parallel, drawing on the aggregate bandwidth of the local DRAM plus the CXL fabric. Figure 8(c) shows such a layout with two AICs. By distributing traffic across multiple cards, the strategy capitalizes on CXL’s aggregate bandwidth, improves scalability in GPU transfer phases, and alleviates memory‑capacity pressure during CPU‑bound steps.

### v2

@@ II-B Memory Bottleneck under Long-Context Offloading
Figure 5: Latency of the CPU-based Adam optimizer step with a growing number of parameters.
Figure 6: Comparison of GPU transfer bandwidth from system DRAM and CXL Memory.
(a) Single CXL AIC
(b) Dual CXL AICs
Figure 7: Latency breakdown of CPU offloading: local DRAM baseline vs. naive CXL configurations.
@@ II-C CXL-Attached Memory
For CXL-attached memory, Figure 4 illustrates the differences in data paths and latency between local and Type 3 devices. Accessing local memory follows a direct path from the CPU cores through the CPU cache and memory controller, resulting in latencies between 80 and 140 nanoseconds [23]. In contrast, accessing CXL-attached memory involves traversing the PCIe interface using the CXL.io and CXL.mem protocols. This path requires coordination between the CPU and the CXL memory controllers, leading to increased latency ranging from 170 to 250 nanoseconds [23]. While CXL-attached memory is utilized as a system memory extension, the Linux kernel integrates CXL-attached memory as CPU-less Non-Uniform Memory Access (NUMA) nodes [33]. This integration allows CXL-attached memory to be managed alongside traditional DRAM, while still enabling users to control allocation explicitly, such as through numactl or libnuma [35], to direct specific data to DRAM, CXL memory, or an interleaved round-robin fashion among available NUMA nodes [35]. Although CXL integration expands usable system memory, our analysis shows that CPU-offloaded long-context LLM fine-tuning remains suboptimal when backed by CXL memory under current deep-learning frameworks (see Section III).
III Observation and Analysis
This chapter empirically characterizes the performance impact of storing CPU-offloaded data on CXL-attached memory. It begins by identifying a critical limitation in current deep learning frameworks that prevents fine-grained memory control. Subsequently, it analyzes the performance of distinct workload components, such as CPU-based computations and GPU data transfers, when using CXL memory. The analysis reveals that naive CXL integration leads to significant end-to-end performance degradation. These findings collectively establish the motivation for the CXL-aware data placement strategies detailed later in this study.
III-A PyTorch Memory Allocation Limitation
PyTorch employs a layered execution stack that lowers high-level Python operations to efficient C++ back-end routines responsible for memory allocation and data movement. This architecture works well in homogeneous memory systems; however, it exposes a central limitation in heterogeneous settings with CXL-attached memory. That is, the current memory placement is controlled at the process level, not at the granularity of individual tensors. In current deployments, NUMA policies configured externally (e.g., via numactl) are applied uniformly to the entire Python process. For example, launching a script with numactl --interleave=0,1 causes all tensor allocations in that process to be interleaved across NUMA nodes 0 and 1, regardless of their role or access pattern. Because PyTorch provides no native mechanism to annotate tensors with placement hints or to route specific allocations to distinct memory tiers, users cannot express policies such as “pin latency-critical optimizer states in local DRAM” while “spilling bandwidth-tolerant or cold tensors (e.g., checkpointed activations) to higher-capacity CXL memory.”
This mismatch is particularly limiting for CPU-offloading pipelines, where different tensor classes exhibit markedly different sensitivity to latency and bandwidth. A single, process-wide policy forces heterogeneous data into a uniform treatment, obscuring opportunities to reduce stall time on the critical path while leveraging capacity elsewhere. The result is suboptimal end-to-end performance and wasted hardware flexibility in mixed-memory environments. These observations motivate a fine-grained allocation extension that (i) exposes per-tensor placement control, (ii) integrates with existing PyTorch allocators without disrupting user code, and (iii) enables principled policies that align tensor characteristics with the most suitable memory tier.
III-B CPU Offloading Slowdowns on CXL-Attached Memory
@@
As illustrated in Section II-A, the CPU‑offloading workflow stores full‑precision parameters, gradients, and optimizer states in system memory so that the CPU can perform the optimizer update locally. As each parameter update is independent, the optimizer phase exhibits ample parallelism. Practical implementations, such as ZeRO-Offload, exploit OpenMP threads and SIMD instructions (e.g., AVX2) to accelerate this compute-intensive step [16]. Consequently, the optimizer phase is highly parallelized and sensitive to increased latency when accessing offloaded data structures. To quantify how memory placement affects optimizer latency within the long‑context CPU‑offloading workflow, the CPU‑based Adam optimizer is benchmarked with offloaded data structures residing either in local DRAM or in CXL-attached memory. Figure 7 summarizes the results. For each configuration, the data size is varied to emulate different LLM scales. Hardware details are listed in Table II, Config. A.
@@
In Figure 7, an “element” consists of a 4-byte parameter, a 4-byte gradient, and 8 bytes of optimizer state. The SIMD kernel processes each element in three steps: (i) it loads the parameter, gradient, and state from memory into vector registers; (ii) executes the floating‑point update; and (iii) writes the updated values back. For modest data volumes, the latency penalty of CXL‑attached memory is negligible; however, once the element count exceeds roughly 20 million, optimizer time with CXL-attached memory rises sharply, reaching nearly 4 times the DRAM baseline. The primary cause is the higher access latency of the CXL path (170-250 ns) compared to local DRAM (80-140 ns), as shown earlier in Figure 4. These results indicate that naively placing latency‑critical optimizer data in CXL‑attached memory can severely degrade fine‑tuning performance. Effective CXL deployments for long‑context CPU‑offloading need to keep latency‑sensitive data in low‑latency DRAM and relegate latency‑tolerant data to CXL‑attached memory, respectively.
III-C GPU Data Transfers on CXL-Attached Memory
@@
The CPU-offloading workflow involves not only a CPU-intensive optimizer step but also frequent, high-volume data transfers between system and GPU memory. For example, before each layer’s backward computation, model parameters and checkpointed activations are copied from system memory to the GPU. Afterward, the resulting gradients are copied back. To quantify the impact of physical memory location on data transfer behavior, this study evaluates GPU copy performance with source buffers placed in either local DRAM or CXL-attached memory. In each experiment, page-aligned host buffers are allocated and pinned to a specific NUMA node. These buffers are registered to enable direct DMA transfers over PCIe, thereby bypassing intermediate copies through CPU caches. Asynchronous memory transfers are then issued, and the resulting effective bandwidth is measured. Figure 7 shows the results.
Figure 8: System architecture of the proposed methods, in which a fine-grained allocation extension adds per-tensor placement control and a CXL-aware memory allocator directs tensors to local DRAM or CXL-attached memory within the existing software stack.
As shown in the figure, the observed transfer bandwidth from CXL-attached memory closely matches that of local DRAM. This is attributed to the use of direct DMA over PCIe, which allows CXL memory to bypass the CPU and transfer directly to the GPU. Since local DRAM also relies on PCIe for such transfers, both memory types share a similar topology. Throughput increases with transfer size until it saturates the PCIe interface bandwidth. This convergence occurs because page-locked buffers expose equivalent DMA paths across both memory types, making the operation bound by interface limits. Furthermore, GPU data transfers tend to tolerate latency more effectively. This is because, for example, CPU-offloading systems can utilize prefetching or asynchronous offloading to mask latency. These results indicate that for high-throughput, latency-tolerant operations such as GPU transfers, CXL-attached memory can deliver performance comparable to local DRAM. This contrasts with latency-sensitive phases, such as optimizer steps, which reinforce the importance of workload-aware data placement.
III-D End-to-End Fine-Tuning Slowdown
@@
Based on the characteristics outlined in previous sections, this section measures how naively incorporating CXL-attached memory affects end-to-end performance during LLM fine-tuning. The comparison evaluates a local DRAM baseline against configurations that combine local DRAM with CXL memory under a naive interleaving policy. This naive approach is representative of what occurs due to the PyTorch memory allocation limitations described in Section III-A. Figure 7 presents the latency profile for fine-tuning a 12-billion-parameter model. In Figure 7(a), a single CXL AIC is used. The optimizer step, labeled as the STEP phase, suffers the most significant slowdown. This is because its CPU-bound loads and stores are acutely sensitive to the higher access latency of CXL memory, a direct consequence of the naive interleaving policy placing critical optimizer data on the slower tier. Phases dominated by GPU transfers, specifically FWD and BWD, exhibit smaller slowdowns because prefetching and asynchronous DMA obscure some of the added latency.
Figure 7(b) shows the result of adding a second CXL AIC. With more available bandwidth from the additional device, the overall performance improves, and the slowdown is mitigated compared to the single-AIC case. However, a performance gap still remains relative to the DRAM-only baseline. This demonstrates that while adding more hardware can alleviate bandwidth issues, it does not resolve the fundamental problem of latency sensitivity in the optimizer step. The naive, process-level memory policy remains a bottleneck. These findings underscore the need for a more intelligent, CXL-aware data placement strategy. Such a strategy needs to distinguish between latency-sensitive and latency-tolerant data and leverage the aggregate bandwidth of multiple CXL AICs effectively.
@@ IV CXL-Aware Long-Context LLM Fine-Tuning
To address the performance degradation arising from the naive adoption of CXL-attached memory, this work introduces a two-part methodology that overcomes the limitations of existing deep learning frameworks and intelligently manages data placement in heterogeneous memory systems. The first component is a fine-grained, per-tensor memory allocation extension for PyTorch, providing essential control over NUMA memory policies. The second is a CXL-aware memory allocator that leverages this extension to strategically place tensors based on their latency sensitivity. Figure 8 illustrates the integration of these components into the LLM fine-tuning stack, where the extension operates between Python and C++ layers of PyTorch, and the CXL-aware allocator directs memory decisions from a higher level.
IV-A Fine-grained Memory Allocation Extension
As discussed previously, on the path to exploiting CXL-attached memory under the CPU-offloading scenario, a significant limitation of existing deep learning frameworks, such as PyTorch, lies in their coarse-grained memory management. By default, NUMA policies are enforced at the process level, causing all tensors allocated within a script to share the same placement strategy. This design prevents applications from selectively assigning tensors to different memory tiers according to their individual access patterns and latency sensitivities. To address this limitation, a fine-grained memory allocation extension for PyTorch is developed to introduce per-tensor control over NUMA placement, allowing developers to allocate tensors directly to local DRAM, a designated CXL device, or an interleaved set of nodes. The implementation is shown in Figure 9. It consists of a Python wrapper that interfaces with C++ backend functions and uses the libnuma library, specifically, numa_alloc_onnode for single-node placement and numa_alloc_interleaved_subset for customized interleaving across nodes. To support efficient GPU data transfers, the extension leverages cudaHostRegister from the CUDA toolkit to pin host memory, thereby enabling zero-copy direct memory access (DMA), while torch.from_blob is used to construct a PyTorch tensor that references the externally managed memory block without additional data copies. This extension establishes the foundational mechanism required for the CXL-aware memory allocator and enables future exploration of advanced memory management strategies in heterogeneous systems.
Figure 9: Workflow of the fine-grained memory allocation extension for PyTorch.
IV-B CXL-Aware Memory Allocator
To effectively exploit the aforementioned fine-grained, per-tensor control for the PyTorch memory allocation extension, this study further develops a CXL-aware memory allocator, which is a runtime algorithm that dynamically partitions CPU-offloaded data between local DRAM and CXL-attached memory. The allocator is designed to minimize performance degradation caused by CXL’s higher access latency by prioritizing data placement based on each component’s latency sensitivity. Its design builds on two key ideas: (1) hierarchical grouping of data by latency tolerance, (2) a latency-first greedy allocation strategy that enforces this hierarchy. In Section IV-B3, the adaptive placement behavior is demonstrated under varying hardware configurations and memory pressure to demonstrate its effectiveness. Together, these mechanisms enable efficient and transparent utilization of heterogeneous memory resources during long-context LLM fine-tuning.
IV-B1 Latency Sensitivity Classification
The allocator classifies CPU-resident data into latency-tolerance levels based on access patterns observed during fine-tuning. Comparing CPU and GPU access behaviors shows that data both accessed and computed on the CPU lies on the critical path and is therefore highly latency-sensitive. For GPU accesses under CPU-offloaded training, this study observes two behaviors: fetch (CPU→GPU transfers required before a layer’s computation) and offload (GPU→CPU transfers performed after computation completes). Because training proceeds layer by layer, fetches must complete immediately (e.g., prefetching bf16 weights for the next forward or backward layer), whereas offloads can be asynchronous (e.g., checkpointed activations after the forward pass, or gradients after the backward pass). Based on the fetch and offload behaviors, three conditions are defined, which are $C_{1}$ compute on the CPU, $C_{2}$ fetching by the GPU, and $C_{3}$ offloading by the GPU. These conditions induce four latency-tolerance levels used by the allocator.
  • Level 1 (Lowest tolerance). Allocations satisfying $C_{1}$ (tight load–compute–store loops on the CPU), such as optimizer states, master weights, and master gradients.
  • Level 2 (Low tolerance). Allocations satisfying only $C_{2}$ (GPU fetch), for example, bf16 weights prefetched prior to a layer’s forward or backward pass.
  • Level 3 (Medium tolerance). Allocations satisfying both $C_{2}$ and $C_{3}$, where accesses are partly hidden by prefetch and asynchronous offload; for example, checkpointed activations offloaded after the forward pass and later fetched for recomputation during backpropagation.
  • Level 4 (Highest tolerance). Allocations satisfying only $C_{3}$ (purely asynchronous offload) that do not lie on the immediate critical path, such as per-layer gradients transferred to CPU memory after computation.
Input: $S_{\text{local}}$: Local DRAM size;
$S_{\text{cxl}}$: Aggregated CXL size;
$num_{\text{cxl}}$: Number of CXL devices;
group_items_sizes: maps each latency level to its list of item sizes $\{level:[size_{1},\dots]\}$
Output: allocations: item $\rightarrow$ (policy, interleave ratio)
$S_{\text{remain}}\leftarrow S_{\text{local}}$;
allocations $\leftarrow\{\}$;
for $level\leftarrow 1$ to 44 do
 foreach $size_{i}$ in group_items_sizes[level]group\_items\_sizes[level] do
    $key\leftarrow$ "level_levellevel_item_ii";
    if $S_{\text{remain}}\geq size_{i}$ then
       allocations[keykey] $\leftarrow$ (pure_local, 1:01{:}0)
       $S_{\text{remain}}\leftarrow S_{\text{remain}}-size_{i}$;
    else if $S_{\text{remain}}>0$ then
       allocations[keykey] $\leftarrow$ (local_cxl, $(1{:}\underbrace{1{:}\dots:1}_{num_{\text{cxl}}})$);
       $S_{\text{remain}}\leftarrow 0$;
    else
       allocations[keykey] $\leftarrow$ (pure_cxl, $\underbrace{1{:}\dots:1}_{num_{\text{cxl}}}$);
return allocations;
Algorithm 1 Latency-First Allocation
Figure 10: Allocation example for single CXL device and sufficient DRAM.
Figure 11: Allocation example for multiple CXL devices and sufficient DRAM.
Figure 12: Allocation example for multiple CXL devices and limited DRAM.
IV-B2 Latency-First Greedy Algorithm
Building upon the four-level latency classification, the latency-first greedy algorithm determines how CPU-resident data are distributed across DRAM and multiple CXL-attached memory devices. The guiding principle of the algorithm is to prioritize low-latency data placement in local DRAM while progressively offloading or interleaving less latency-insensitive data into CXL memory as DRAM capacity becomes limited. This design follows a greedy strategy, which makes locally optimal decisions at each step, so that latency-critical data always receive preferential treatment without incurring the complexity of global optimization or iterative tuning. Algorithm 1 presents the detailed allocation process. The algorithm takes as input the available local DRAM capacity ($S_{local}$), the aggregated capacity of all CXL devices ($S_{cxl}$), and the number of CXL devices ($num_{cxl}$). The CPU-resident data are grouped by latency sensitivity into the mapping structure group_items_sizes, which associates each of the four tolerance levels with a list of tensor or parameter sizes. The output is an allocation table that maps each data item to a specific placement policy and its corresponding interleave ratio across memory tiers.
The allocator begins with the full DRAM capacity ($S_{remain}$$S_{local}$) and iterates through the latency levels from the most sensitive classification (Level 1) to the least sensitive classification (Level 4). Within each level, every item is evaluated based on its size and the remaining DRAM space. If an item completely fits within the available DRAM ($S_{remain}\geq Size_{i}$), it is placed entirely in local DRAM, marked as pure_local, ensuring minimal latency for the most performance-critical components such as optimizer states and master gradients. If DRAM space is insufficient but non-zero ($S_{remain}>0$), the allocator interleaves the data across the remaining DRAM and all CXL devices, adopting the local_cxl policy with an even interleave ratio. This configuration allows the DRAM portion to serve as a fast buffer while using CXL memory to absorb the overflow and take advantage of the bandwidth from multiple CXL memory devices. Once DRAM is exhausted ($S_{remain}=0$), the remaining items are distributed evenly across all CXL devices using the pure_cxl policy, thereby maximizing aggregate bandwidth and overall capacity.
This greedy traversal ensures that DRAM is always reserved for the most latency-critical tensors while maintaining balanced utilization of the heterogeneous memory system. The resulting allocation map explicitly specifies both the placement policy and interleaving configuration, allowing the runtime system to reproduce consistent memory behavior across runs. By following a latency-first decision order, the allocator achieves an effective compromise between minimizing response time for critical data accesses and exploiting the extended capacity of CXL memory. Furthermore, its deterministic, lightweight design makes it suitable for integration into existing deep-learning runtimes without requiring online profiling or costly dynamic migration during fine-tuning.
IV-B3 Allocation Examples
The behavior of the CXL-aware allocator is illustrated through three representative scenarios that reveal how the algorithm adapts to different hardware configurations and memory pressures.
  • Scenario 1 (Single CXL Device, Ample DRAM). As shown in Figure 12, a system equipped with one CXL device and sufficient DRAM places all latency-critical Level 1 data, which includes optimizer states (fp32 O), master parameters (fp32 P), and master gradients (fp32 G), entirely in DRAM. The more latency-tolerant data, such as checkpointed activations (bf16 A), parameters (bf16 P), and gradients (bf16 G), are stored in the CXL memory.
  • Scenario 2 (Multiple CXL Devices, Sufficient DRAM). In Figure 12, multiple CXL devices are introduced while DRAM capacity remains sufficient for Level 1 data. Latency-sensitive data stay in DRAM, whereas latency-tolerant data (bf16 A, bf16 P, bf16 G) are interleaved across the CXL devices to exploit their aggregate bandwidth, enhancing throughput for GPU data transfers.
  • Scenario 3 (Multiple CXL Devices, Limited DRAM). As illustrated in Figure 12, when DRAM is insufficient to accommodate all Level 1 data, the greedy allocator first fills the available DRAM with the highest-priority tensors. The remaining Level 1 data, such as part of the optimizer states (fp32 O), are striped across the remaining DRAM and all CXL devices. This fallback mechanism ensures balanced utilization of memory resources while minimizing the performance penalties associated with higher-latency CXL access.
(a) 7B model in a single-GPU scenario
(b) 7B model in a dual-GPU scenario
(c) 12B models in a single-GPU scenario
(d) 12B models in a dual-GPU scenario
Figure 13: Training throughput comparison of three single-AIC configurations: Baseline (local DRAM only), Naive CXL (interleaving), and CXL-Aware Allocation (labeled as Our), evaluated across varying models, context lengths and batch sizes.